What will happen to the current with 5 VRMS and 0.22F connected in series as frequency increases?
1
Expert's answer
2020-09-09T10:21:02-0400

Explanations & Calculations


  • Capacitors, like resistors show some resistance—normally called as the impedance (Z) of the capacitance (C)— when used in ac circuits & it's given by, "\small Z = \large \frac{1}{C\omega}"
  • Then,

"\qquad\qquad
\begin{aligned}
\small Z &= \small \frac{1}{C\times 2\pi f}
\end{aligned}" where "\small f" is the frequency of the supply implies that as the frequency increases the value of Z decreases.

  • According to the generalized Ohm's law, "\small V = i Z \to i=\large \frac{V}{Z}" , "\small i" is expected to increase as the applied voltage remains constant.


  • For this sum C = 2.2F & Vrms= 5V